MIDC News
Year of 2002
December 17, 2002
- 14:00 MST: All BMS Radiometer Tower instruments now are working. The BMS QA Pages for the following instruments has been updated to indicate the data quality: Global Normal CM21, Global 40-S PSP, Global 40-S LI-200, Upwelling PIR, and Upwelling PSP. -Afshin Andreas
- 12:46 MST: Radiometer tower j-box replacement complete as of 12:30 today. Li-200 still requires change in sign of calibration coefficient. -Pete Gotseff
- 08:10 MST: The J-Box on the SRRL BMS radiometer tower was removed for replacement yesterday at approx 14:30 MST. All instruments on the radiometer tower (except the CNR1) were affected. Most instruments are still reading bogus values at this time. -Afshin Andreas
- 12:41 MST: The Wind Chill Calculation for both SRRL BMS and NWTC M2 was updated today at 11:57 MST (BMS) and at 12:20 MST (M2). We are now using the new National Weather Service (NWS) method as described on the NOAA web site. Besides an updated formula, the new method uses wind speed at 5 feet rather than 33 feet (old method). Since neither BMS or M2 has wind sensors at exactly 5 feet, the following was used:
- BMS: 10 feet (old method was at 42 feet)
- M2: 2 meters [6.6 ft] (old method was at 10 meters [33 ft]).
- Wind Speed > 1.34112 m/s (3 mph)
- Temp < 20oC (68oF)
- Wind Chill < Temp
- 15:44 MST: All the SRRL BMS Monthly Data Sets for November 2002 have been generated. -Afshin Andreas
- 15:19 MST: A bug in the report generator for SRRL BMS was found, which affects instruments that are supposed to have daily averages during daylight hours. This problem will be fixed for future reports. The October 2002 and prior reports for the (13) affected instruments will show a lower value for the daily average (due to the zero night time hourly values being averaged in). The (13) affected instruments include the following: Global UV Index, (5) Turbidities, (2) AODs, (2) Albedos, Air Mass, and Zenith & Azimuth Angles. -Afshin Andreas
- 08:26 MST: The datalogger program for SPMD was updated today @ 8:05MST to include 4 additional outputs: Cabin Temp, Cabin Lower Temp, CR10 Battery Voltage, and 24V Battery Voltage. Historical data since Sept 6, 2002 was reformated to allow for the change. The (2) temps were set to -99999 and the (2) voltages were set to the sampled value @ midnight for the entire day for the historical data. -Afshin Andreas
- 13:16 MST: The SRRL RSP was experiencing intermittant problems with the pyranometer temperature reading. On 11/19/02 at 15:30 hours the connections for the temperature sensor were checked and tightened. Temperatures look ok for now.
-Pete Gotseff
- 14:35 MST: Output limits were placed on the following MIDC products:
- SkyCam Gallery: Max. 1000 Thumbnails
- SkyScanner Data: Max. 1 Year Span
- Spectral comma-delimited data: Max. 1 Year Span
- Spectral two column data: Max. 1 Month Span
- 11:46 MST: A MIDC Search page is now available, powered by the NREL search engine. -Afshin Andreas
- 11:01 MST: Updated the NWTC M2 Instruments page to include calculated fields. -Afshin Andreas
- 08:51 MST: Links to manufacture instrument information has been added to the SRRL BMS Instrument page by clicking on the model numbers. -Afshin Andreas
- 12:11 MST: Minor change was made to DQMS QA Testing for SRRL BMS Global UVB instruments. -Afshin Andreas
- 12:01 MST: SRRL BMS Monthly Data Sets/Reports for October 2002 have been generated. -Afshin Andreas
- 13:30 MST: The SRRL EKO Sky Scanner stopped functioning on the morning of October 29. Problem is unknown at this time. -Afshin Andreas
- 13:26 MST: The BMS Global 40-South LI-200 was reading low, the instrument was replaced with a s/n PY1724 (calibrated 9-2-02) on Oct 10, 2002 @ 12:48 MST. -Afshin Andreas
- 09:01 MST: The SPMD windrose page (monthly, seasonal, & yearly) has been activated. Currently one rose for the month of October 2002 is available. -Afshin Andreas
- 12:32 MST: The SRRL BMS Instrument Page has been updated to include calibration links for most of the remainder of non-Infrared and non-Meterological instruments. -Afshin Andreas
- 12:44 MST: The SRRL BMS Instrument Page has been updated, to include links to our AIM Database that will display a calibration history for the instrument serial number that has been clicked on. -Afshin Andreas
- 08:10 MST: The Lamar Low-Level Jet Project (LLLJP) home page is now visible to all users, however the actual data can only be downloaded by NREL users using the old internal URL. This site has been on the MIDC since Oct 2, 2001. -Afshin Andreas
- 07:38 MST: External (non-NREL) Users must now use the new URL of http://www.nrel.gov/midc/ to access the MIDC. Internal NREL users can use the new URL or the old URL (http://midc.nrel.gov), however the old address is recommended for faster perfomance. -Afshin Andreas
- 09:24 MST: The NWTC M2 historical Accumulated Global data (8/24/2001 to 9/10/2002), was corrected by doubling the value as stated by Sept 11, 2002 news item. -Afshin Andreas
- 09:22 MST: The NWTC M2 historical aspirator old signal data (8/24/2001 to 1/24/2002) was corrected to reduce confusion. If the signal was approx 5 VDC (bad), the value was set to 0 RPM, otherwise it's 3000 RPM. -Afshin Andreas
- 09:20 MST: The NWTC M2 historical data since the CR23x was installed has been corrected (8/24/2001 @ 16:05 to 9/26/2002 @ 09:48). The following RTDs were corrected:
- DewPoint @ 2m was increased by 1.4 oC
- Temp @ 2m was increased by 1.4 oC
- Temp @ 50m was increased by 2.8 oC
- Temp @ 80m was increased by 3.7 oC
After correcting the RTDs, all affected calculations were re-calculated, which includes:- Wind Chill Temp
- Relative Humidity
- Specific Humidity
- Richardson Number (2-50m)
- Richardson Number (2-80m)
- Richardson Number (50-80m)
The CR23x is now reading a more accurate temperature for the RTDs as stated by Sept 27, 2002 news item. -Afshin Andreas - DewPoint @ 2m was increased by 1.4 oC
- 13:38 MST: Added a new feature to the SRRL Spectral Plot page: There is now a download link for each instrument displayed on the plot, that allows the user to download the data for the particular time shown on the plot.
-Afshin Andreas - 08:04 MST: The SRRL Direct LI-1800 was returned to service today @ 7:55 MST. -Afshin Andreas
- 15:25 MST: The PIR and CG4 had their shading disks removed from 1300-1315 hours today for brief testing. -Pete Gotseff
- 10:59 MST: The BMS Direct LI-1800 fiber probe has been removed for service. Water and dirt leaked into the lens housing of the fiber probe which is attenuating the signal. -Afshin Andreas
- 07:20 MST: The Measurement of the BMS CNR1 Case Temperature was upgraded on Oct 7 @ 13:06, the logger now measures the PT-100 using Instruction 9 (4 wire half bridge) instead of Instruction 8 (differential meas. w/excitation). The temperature using the old method was approx. 0.5 degrees low. This temperature is also used in deriving the Upwelling & Downwelling IR CG3s. -Afshin Andreas
- 07:16 MST: The SRRL LI-1800 Direct Normal Data is low by about 3, problem is unknown at this time. Last good data point was on Oct 1 @ 11:00 MST. -Afshin Andreas
- 15:30 MST: All SRRL BMS Monthly Data Sets & Reports for September 2002 have been generated. -Afshin Andreas
- 08:46 MST: The SPMD Pictures have been updated to show the new Met One Sensors. -Afshin Andreas
- 09:34 MST: The correct cal factor for the SPMD SS-201 was entered into the logger today @ 9:35 MST. -Afshin Andreas
- 06:55 MST: The Texas Electronics TV-102 Wind Speed sensor at SPMD was replaced with a Met One SS-201 on Sept 28 @ 13:55 MST. The correct calibration factor has NOT been entered into the logger yet. -Afshin Andreas
- 11:14 MST: The SRRL EKO SkyScanner is now functional as of 10:00 MST. Source of problem is unknown. -Afshin Andreas
- 11:12 MST: The NWTC M2 Temperature and Dew Point Sensor measurements were upgraded yesterday @ 10:00MST. The logger now measures the PRTs using Instruction 9 (4 wire half bridge) instead of Instruction 8 (differential meas. w/excitation). The temperatures with the old method were approx. 1.5 degrees C to low. The new method has corrected this problem. -Afshin Andreas
- 07:17 MST: The SRRL EKO SkyScanner stopped functioning on Sept 20 @ 16:15 MST. Problem is unknown at this time. -Afshin Andreas
- 09:39 MST: All SRRL BMS Monthly Data Sets for August 2000 through October 2000 has been generated. -Afshin Andreas
- 13:48 MST: Fixed minor bug in solar calendars which caused thumbnail plots to show positive missing values (99999). -Afshin Andreas
- 13:45 MST: Fixed minor bug in solar calendars which caused thumbnail plots to autoscale (instead of being fixed at 1200) when there is missing data on the first day of the month. (example BMS Nov 1985) -Afshin Andreas
- 13:44 MST: Repaired 8 BMS daily data files in 1987 to 1993 that had minor data corruption. -Afshin Andreas
- 11:25 MST: SRRL BMS QA Pages for Global PSP, Direct NIP, and Diffuse PSP, were updated describing what occured on Aug 2000 through Oct 2000. -Afshin Andreas
- 07:51 MST: The SPMD Wind Roses have been corrected....The Wind Speed data is converted from MPH to m/s on the rose. -Afshin Andreas
- 06:48 MST: The new wind direction sensor was installed at SPMD on Sept 14 @ 13:30 MST. Note: The units on the SPMD wind rose is wrong (it should be MPH). This will be corrected soon. -Afshin Andreas
- 15:03 MST: SRRL BMS Daily data files for August 2000 through October 2000 are now available. SRRL BMS was undergoing upgrade and reinstallation on the new instrument deck at that time, however SRRL BSRN was still operational. The Data needed some correction before making part of BMS. The Direct NIP data was good, the Global PSP data was increased by 8%, 25 days out of the 3 months the Diffuse was calculated based on global and direct because the tracker was malfunctioning. Also 4 days both the direct and diffuse trackers were out, the data for for those instruments (on those days) were set to missing values. These days will be listed on the BMS Instrument QA page shortly. Monthly data for these 3 months will be generated sometime soon. -Afshin Andreas
- 09:40 MST: All BMS Monthly Data for August 2002 has been generated, which include Monthly Data/QA and Monthly Reports/Plots. -Afshin Andreas
- 07:07 MST: The calculation of the M2 Global Accumulated Irradiance was corrected today @ 7:02 MST. The calculation was off by a factor of 2. This problem was caused by the datalogger execution interval being increased from 1 to 2 seconds). Historical data will be corrected soon. -Afshin Andreas
- 17:10 MST: A new data format was created for SPMD starting today, to allow for a wind direction instrument that will be added to the site tomorrow. -Afshin Andreas
- 11:18 MST: Another plot was added to the BMS instrument page for the CIE 1987 Erythema Action Spectrum and the 501A Response. -Afshin Andreas
- 10:54 MST: Two plots are now available on the SRRL BMS Instrument Page for the UV Instruments, which show the UVB Spectral Response and the UVA/Total Spectral Response for each instrument. -Afshin Andreas
- 15:09 MST: A BMS Instrument QA page was created for the Diffuse CM22 listing the date & times, when the shading ball was removed during calibration. -Afshin Andreas
- 08:51 MST: The cal factor for the BMS Diffuse CM22 and Direct LI-201 were updated in the logger today @ 8:48 MST. (They were calibrated in BORCAL 2002-04) -Afshin Andreas
- 07:38 MST: The old tower global LI-200 sensor at SPMD was removed on Aug 31 @ 11:00 MST. Historical data from 8/6/2002 through 8/31/2002 for the tower sensor was removed and replaced with the new roof sensor data. -Afshin Andreas
- 08:24 MST: Minor documentation error was discovered on SRRL BMS Reports (November 2000 - July 2002): Licor LI-190 has wrong units in title and calibration factor; Kipp & Zonen CUVB2 has wrong serial number. Error will be fixed for future reports. -Afshin Andreas
- 14:32 MST: BMS Instrument History page has been updated to show new cal factors for all UV Instruments. -Afshin Andreas
- 15:18 MST: The cal factors for the BMS UV-Index and UV-MED/Hr were updated in the logger today @ 15:16 MST. -Afshin Andreas
- 14:56 MST: The BMS UV Instruments were calibrated and the new factors were entered into the logger at 14:48 MST: Global TUVR, UVSAT, CUVA1, UVSBT, CUVB1, MS210W, UVB1, 501A; Direct TUVR, CUVA2, CUVB2. -Afshin Andreas
- 09:06 MST: The Shading Ball was removed on the SRRL BMS Diffuse CM22 on Aug 24, to allow it to be calibrated in a BORCAL. -Afshin Andreas
- 17:14 MST: The SRRL Global 40-South LI-1800 was returned to service today @ 14:40 MST. The SRRL Direct Normal LI-1800 was returned to service today @ 15:25 MST. -Afshin Andreas
- 10:54 MST: The (2) SRRL LI-1800 Spectroradiometers were removed at 8:00 MST today for calibration. -Afshin Andreas
- 09:14 MST: Troubleshoooting revealed that the EKO SkyScanner head cable was defective. Cable has been replaced with a good used one. New one is on order. -Pete Gotseff
- 07:04 MST: The EKO MS-300LR SkyScanner failed on Aug 4, 2002, both the luminance and radiance measurements are bad. -Afshin Andreas
- 06:59 MST: Another Global LI-200 pyranometer is going to be added to SPMD, later today. The current LI-200, which is mounted on the tower, is being shaded in the morning. The new LI-200 is going to be mounted on the roof. Both the new and old LI-200 will remain installed for about a month to allow for comparison. A new "data format" was created today to allow for this change. -Afshin Andreas
- 07:17 MST: Yesterday it was found the back of the Direct RG780 NIP filter had accumulated an excessive amount of dirt. Cleaning it resulted in an approx. 25W/m2 increase in irradiance values ... it is unclear how long this problem had existed.
-Pete Gotseff
- 16:35 MST: A page has been created to describe the UpWelling ShortWave Test that is used to test BMS Instruments (7/2002 to present). -Afshin Andreas
- 15:14 MST: SRRL BMS July 2002 Monthly Data has been processed using DQMS 3, the NREL Cylinder/QCFit Plots, and the MIDC Report Generator/Flag Summary. All products for July are available now available on the BMS web site. -Afshin Andreas
- 14:06 MST: The BMS Downwelling PIR was swapped with a reference PIR on 7-30-02 @ 10:08 MST through 8-1-02 @ 13:51, to test the performance of the new NREL/SGP BB Calibration Coefficients. -Afshin Andreas
- 12:39 MST: Added the option of plotting instrument vs day on the BMS Monthly Reports/Plots page. -Afshin Andreas
- 10:15 MST: The SRRL BMS Monthly Plots (derived from the reports) showing the minimum, average, and maximum values over the month for each hour of the day are now available for all historical data (any instrument) on the Monthly Reports and Plots page. -Afshin Andreas
- 07:02 MST: All BMS CSV ASCII Reports have been generated again, using a new
limit of 90% (instead of 80%). The PDF files have also been regenerated to reflect this change and the two changes mentioned yesterday. -Afshin Andreas
- 06:51 MST: Two problems were found in the new BMS Reports: 1. The header line should read "W Longitude". 2. Some February Reports had a day 29 that did not belong (all values were zero for the day). All the reports have been reproccessed and the CSV ASCII versions have been corrected. The corrected PDF versions will be available to tomorrow. -Afshin Andreas
- 09:06 MST: All SRRL BMS Historical Reports are now up to date, this includes April 1985 through June 2002. Old reports that were generated from the VAX (QC_TND) and DQMS 2 have been re-generated using the new report generator. This is a total of 5,173 instrument reports for the 204 months. The Reports are now available as both CSV ASCII files and PDF Files.
-Afshin Andreas
- 09:24 MST: The following SRRL-BMS Instruments were calibrated in BORCAL 2002-02 (on 6-6-2002): Diffuse CM22, Direct NIP RG780, Global PSP, Direct NIP 1 & 2, Diffuse 8-48. The new cal factors were entered into the logger program today @ 7:32 MST. The BMS Instrument page has been updated. -Afshin Andreas
- 09:18 MST: The SRRL-BMS-Deck Datalogger program was updated today @7:50 MST, to be compatible with the new NREL/SGP Black Body Calibration Coefficients. This affects the following instruments: Downwelling PIR, Upwelling PIR, and Downwelling CG4. The BMS Instrument page has been updated to show the new calculation. -Afshin Andreas
- 10:17 MST: BMS QA pages were created to indicate a past problem with the Global 40-South PSP and the Direct RG780 NIP. -Afshin Andreas
- 10:53 MST: The new BMS Flag Summary Tables for April 1985 to December 1996 have been generated (they never existed as the flag bar chart). -Afshin Andreas
- 10:48 MST: The BMS Flag Summary Tables for Nov 2000 through June 2002 have been generated. -Afshin Andreas
- 10:47 MST: The following changes have been made to the BMS QA'd monthly data: the "Flag Summary Table" is now called the "Full Flag Summary" and the "Flag Bar Charts" have been converted to an ASCII table and are now called the "Flag Summary Table". -Afshin Andreas
- 08:03 MST: The Full Flag Summary for the BMS QA'd Monthly Data from Nov 2000 through June 2002 has been generated. -Afshin Andreas
- 07:13 MST: The BMS Monthly QA'd Data for Nov 2000 through June 2002, has been reprossesed to include zenith, azmuth, and airmass. Also the flags for 34 irradiance instruments, which were tested using a RATIO or MINMAX test, had the flags reset to "00" (untested) during the nighttime hours. -Afshin Andreas
- 11:06 MST: The SRRL BMS QA'd Monthly Data for November 2000 through June 2002 is now available. This includes the quality assessed data, cylinder plots, and qcfit plots. The reports, ghi/dhi/dni plots, and flag summary/plot, are not ready at this time. Note: The new data is also available via the Custom monthly data sets page. -Afshin Andreas
- 08:34 MST: In prepartion for the arrival of the long awaited SRRL BMS QA'd Monthly Data Sets for November 2000 to present, the field defintions page and the QA Tests page has been updated. Also a new QA Flags page and a LongWave Test page has been created. -Afshin Andreas
- 07:02 MST: The BMS Global TUVR overanged the datalogger a couple times on July 5 and 6 (output was -99999), the range on the logger program was increased today from the 10mV to 50mV scale. -Afshin Andreas
- 15:04 MST: Added a Legend to the Solar Calendars Page. -Afshin Andreas
- 13:59 MST: Fixed a problem with the solar calendars (images were not "clickable"), that occurs when a user is using a MAC and Internet Explorer. -Afshin Andreas
- 11:06 MST: Albedo (Eppley) has been added to the SRRL BMS Display. -Afshin Andreas
- 09:18 MST: The option of ZIP Compressing the output from the SRRL BMS Custom Monthly Data is now available. -Afshin Andreas
- 12:40 MST: The SRRL BMS Display has been updated, it now shows Turbidity (TWC 500nm). Wind Sheer has been removed. -Afshin Andreas
- 07:22 MST: The NWTC M2 Global PSP, was replaced with a calibrated instrument on June 28 @ 7:47 MST, the calibration factor was updated in the logger program on July 1 @ 7:09 MST. See the M2 Instrument Page for the current calibration info. -Afshin Andreas
- 12:26 MST: On June 17, 2002 some of the BMS data was reprocessed(see news item for that date below), however during the time frame of Aug 24, 2001 through Sept 13, 2001, the raw photometer output was missing, therefore the turbidity data was not calculated (however this turbidity data was available in the old data file). These data files were corrected today by optaining the turbidity from the old data file and inserting it into the current data files. -Afshin Andreas
- 11:15 MST: A new field for Aerosol Optical Depth was added to BMS 8, to allow for both 500nm AOD and BroadBand AOD. All historcal data since Aug 4, 2001 has been reprossessed to include this change. -Afshin Andreas
- 06:49 MST: An Instrument QA page has been created for the BMS primary NIP to indicate the recent problem. -Afshin Andreas
- 12:04 MST: The problem with the BMS NIP #1, was corrected today @ 11:00 MST. The signal from another instrument (to be calibrated in a BORCAL) was put in parallel with the signal from the NIP, which caused the drop in voltage output. -Afshin Andreas
- 08:07 MST: The calculation for the BMS Aerosol Optical Depth was changed today @ 7:25 MST. It is now the AOD at 500nm instead of BroadBand AOD. Historical data will likely be corrected sometime soon. -Afshin Andreas
- 07:45 MST: The nite-time offset on the BMS Diffuse UVB-1, appears to be fixed. (The instrument was disconnected and reconnected yesterday). -Afshin Andreas
- 07:29 MST: A new field has been added to BMS 8, called "Net Infrared PIRs (atmos.)", all historical data since Aug 4, 2001, has been reprocessed to include this field. -Afshin Andreas
- 14:38 MST: All BMS historical data for the Net Radiation (Eppley), since Aug. 4, 2001 has been corrected. -Afshin Andreas
- 13:00 MST: The calculation of the BMS Net Radiation (Eppley) was corrected today @ 12:43 MST. It was previosly calculated using the net infrared based on the downwelling PIR instrument, it is now calculated based on the actual atmospheric net infrared (downwelling PIR minus upwelling PIR). The historical data will likely be corrected sometime in the near future. -Afshin Andreas
- 12:56 MST: An error with the net infrared naming convention was discovered with BMS. The NET Infrared for the PIR and CG4 is the instrument net (not the atmospheric net, which is equal to downwelling IR minus upwelling IR.) The NET infrared listed for the CNR1 is the actual atmospheric net infrared. The labels have been modified to indicate "(inst.)" and "(atmos.)". A new field may be added in the future to display the atmospheric net infrared based on the Eppley PIRs. -Afshin Andreas
- 10:08 MST: There is a problem with the BMS Direct NIP #1, which started on June 20, it is reading too low. -Afshin Andreas
- 09:47 MST: There is a problem with the BMS Diffuse UVB-1 that occured on June 20, the instrument is showing a large positive nite time offset. -Afshin Andreas
- 10:06 MST: An output limit of 1 Million Values has been placed on SOLPOS, to prevent a user from generating too much data (and consuming too much web server cpu time and temporary disk space). This limit will allow for the following approx. limits (based on one output parameter):
- 11 days of 1 second data
- 694 days of 1 minute data
- 114 years of 1 hour data
- 13:37 MST: When the new SRRL BMS v8.0 was put into service on Aug 4, 2001, not all the calculated fields were in affect (output was -99999). As of today, this data was reprocessed to include all the calculations. The dates affected were 8-4-2001 through 9-14-2001. -Afshin Andreas
- 15:50 MST: Narrow band calibration factors were originally used for the following BMS Kipp & Zonen UV instruments: CUVA1, CUVA2, CUVB1, CUVB2, between 11/1/2000 to 8/3/2001. The data files for all these instruments was updated to use the wide band (full UVA/UVB) calibration factor which was used on 8/4/2001 (which was derived from calibration data taken in 11/2000). The instrument calibration history page has been updated to reflect these changes. -Afshin Andreas
- 15:44 MST: The incorrect calibration factor for the BMS TWC Photometer that was entered from 11/1/2000 to 8/23/2001 was corrected in the datafiles, a correction is no longer required. The TWC QA Page has been updated to reflect this change. -Afshin Andreas
- 09:34 MST: The noise problem with the NWTC M2 Temp @ 50m was corrected on June 11 @ 11:30 MST, by loosening and tightening a screw on the patch panel at the base of the tower. -Afshin Andreas
- 07:52 MST: There is a problem with the SRRL LI-1800 Direct Normal Spectroradiometer, it has been reading very low since June 1. -Afshin Andreas
- 08:12 MST: A temporary calibration factor was derived for the BMS Direct TUVR using the BMS CUVA2 and CUVB2 as the reference, for the range of 280-400nm. The new factor was entered into the logger program today @ 8:08 MST. -Afshin Andreas
- 12:35 MST: At 12:30 MST 6/5/02 the Direct normal TUVR was returned to service after repairs by Eppley Labs. The instrument has not yet been calibrated.
-Pete Gotseff
- 09:39 MST: The BMS Sun-Follower III has been replaced with spare unit (failure possibly due to failed microswitch on AZ drive gear). Returned to service @ 13:00 MST 28 May. -Pete Gotseff
- 08:19 MST: The BMS Sun-Follower III has been down since May 26 @ 00:00 MST, due to tracker failure. The tracker will require replacement or repair. This affects the following direct normal instruments: NIP #1, NIP RG780, CUVA2, CUVB2, LI-201, and TWC 500nm. -Afshin Andreas
- 07:23 MST: As of 7:22 MST today, the BMS Albedo (Eppley) is now calculated using the Ventilated Global PSP (instead of the non-ventilated). -Afshin Andreas
- 06:46 MST: The filter on the BMS Direct NIP RG780 was removed today @ 6:00 MST, to allow the instrument to be calibrated in BORCAL 2002-02. -Afshin Andreas
- 13:38 MST: The EKO Sky Scanner radiance shutter was found to be non-operational this morning. Disassembly of the head unit found the selenoid had become disengaged from the shutter actuator. The Sky Scanner was returned to service at 13:15MST today.
-Pete Gotseff - 13:35 MST: At 13:23MST the Direct TUVR was removed from service. A short has been placed in its place.
-Pete Gotseff - 08:18 MST: The BMS Instrument History Page was updated to include the following changes that occured to the datastream on Sept 13, 2001: -Afshin Andreas
- 07:56 MST: Telecommunication with Elizabeth City State University (ECSU) has been re-established. -Afshin Andreas
- 09:02 MST: A temporary calibration factor was derived for the new BMS Global TUVR (s/n 33444) derived from the BMS Global UV-S-A-T and UV-S-B-T for 280-400nm and entered into the logger at 8:48 MST today. See the instrument history page for more information. -Afshin Andreas
- 07:38 MST: There is telecommunications trouble with Elizabeth City State University, last successfull datacollection occured on May 8 @ 8:35 EST. -Afshin Andreas
- 14:30 MST: As of 14:20 MST today, the SRRL LI-1800 Direct Spectroradiometer is operational. -Afshin Andreas
- 09:29 MST: On May 3 2002, 13:30 The Global TUVR was replaced with a new Eppley TUVR. The calibration factor is not accurate at this time - Global UV is reading high. -Pete Gotseff
- 10:01 MST: Corrected a over-ranging problem with the NWTC M2 Global PSP, which caused an output of -99999, during very sunny conditions (above 1100 W/m2). The datalogger program was updated today @ 9:22 MST. -Afshin Andreas
- 12:05 MST: The SRRL EKO Sky-Scanner was returned to service today @ 11:45 MST. -Afshin Andreas
- 12:40 MST: The SRRL Schott RSP Program was updated to revision 113 today @ 12:36 MST (a correction was made to the direct calculation). -Afshin Andreas
- 09:55 MST: Fixed a bug that caused no data output when generating custom monthly data (this only applies to SRRL BMS). -Afshin Andreas
- 09:36 MST: The GPS Reciever was power cycled and is now working properly. -Afshin Andreas
- 07:36 MST: The GPS Unit at SRRL is malfunctioning, so our time server may be off by a few seconds. -Afshin Andreas
- 06:40 MST: The AOCS is now running as of today @ 6:38 MST. Problem caused by a "Microsoft prompt message", which has been corrected (and won't happen during next power failure). -Afshin Andreas
- 06:39 MST: The Sun-Follower III tracker was reset today @ 6:30 MST. -Afshin Andreas
- 20:50 MST: I forgot to reset the Sun-Follower III tracker. There was no sun today anyway, but there may be sun tomorrow. -Afshin Andreas
- 20:45 MST: The AOCS is also not running for some reason. I did not notice it when I was there at 16:13 MST. -Afshin Andreas
- 16:13 MST: Data is now updated for all sites on the MIDC. -Afshin Andreas
- 14:21 MST: Data on the web server is not updating for any sites on the MIDC today, due to a required power outage at 7:00 MST. Problem caused by a "Microsoft prompt message", which has been corrected (and won't happen during next power failure). -Afshin Andreas
- 10:06 MST: Revision 112 of the SRRL RSP program now outputs pressure/elevation corrected airmass (it was not pressure corrected before). Note: SRRL BMS Airmass is not pressure corrected. -Afshin Andreas
- 09:36 MST: The SRRL RSP Program was updated to revision 112, today @ 8:56 MST. -Afshin Andreas
- 11:15 MST: The BMS Sci-Tec tracker clocks were reset today @ 11:00 MST, they were approx 1 minute off. Before the clocks were reset, a saw-tooth affect could be seen in the MS110 Photometer due to it's very small field of view. -Afshin Andreas
- 09:08 MST: Fixed a bug with the go2url redirector, which caused some links to fail when using a Netscape browser. -Afshin Andreas
- 08:44 MST: I restored the original SCID datalogger program today @ 7:40 PST. -Afshin Andreas
- 08:42 MST: Someone changed the SCID datalogger on April 4, 2002 @ 15:40 PST, without my knowledge. They used a older version of the program which had programming errors (zero output for wind directions and battery voltage). -Afshin Andreas
- 08:34 MST: Power to the BMS Trackers and Direct LI-1800 spectroradiomter was inadvertintly disconnected momentarily at 11:06 MST yesterday. As a result the Direct LI-1800 is still not functioning. -Afshin Andreas
- 07:51 MST: The BMS Diffuse PSP was returned to service yesterday @ 11:40 MST. -Afshin Andreas
- 11:27 MST: The Direct and Global TUVRs have been shorted today at 11:20AM to attempt to reset the detectors which have been reading erratically over the last few weeks.
-Pete Gotseff
- 08:38 MST: There is a problem with the BSRN Diffuse Tracker, which occured yesterday around 9:00 MST. -Afshin Andreas
- 13:11 MST: The BMS Global TUVR is having intermitten problems, which started yesterday around 13:30 MST. -Afshin Andreas
- 09:25 MST: The BMS CG4 was returned to service today at 9:22 MST. -Afshin Andreas
- 14:11 MST: There is a problem with the BMS Direct TUVR again, which started today. -Afshin Andreas
- 12:51 MST: The NWTC M2 Live Display software has been upgraded to allow for both a high res display (1280x1024 for LCDs located in Bldg 251 & IUF) and a low res display (990x740 for web users). This will allow for easier web viewing of the display without scrolling. -Afshin Andreas
- 13:09 MST: The BSRN downwelling diffuse single axis Eppley tracker was replaced 1130-noon 03/06/2002. Previous ST-1 single axis tracker had acquired too much slop in the gears due to high wind. -Pete Gotseff
- 08:16 MST: The BMS Direct TUVR QA Page was updated to indicate the corrected operation of the instrument. A BMS Diffuse PSP QA Page was created to describe the instrument being out of service. -Afshin Andreas
- 07:50 MST: Due to a shortage of diffuse trackers, the BMS Diffuse PSP was removed from service on Feb 26 @ 9:30 MST, to allow for another 8-48 to be calibrated in it's place. The instrument output was shorted on Feb 27 @ 7:47 MST to indicate that it is out of service. Note: This also affects the calculated Diffuse PSP (thermal corrected). -Afshin Andreas
- 08:50 MST: Afternoon Feb 19, 2002 The Direct TUVR sensor was shorted from approx. 1-3pm in an effort to "reset" the detector. The Direct TUVR has been experiencing problems for many weeks with its signal effectively zeroing out. Compare to Direct UVA -Pete Gotseff
- 12:38 MST: The AOCS software has a problem, after a couple days of running the generated parameters are wrong. The software requires a manual restart. A more permanent solution will be implemented when I have free time. -Afshin Andreas
- 10:00 MST: The Direct LI-1800 was returned to service today @ 10:00 MST. It is plugged into UPS power until we get a replacement battery. -Afshin Andreas
- 13:12 MST: Fixed a bug in the NWTC M2 Post-proccessing (today @ 13:03 MST) that affected the calculation of the Richardson Numbers at 2-50m and 50-80m. The bug caused the following to occur: The Richardson No. (50-80m) was calculated based on a constant wind shear of 1e-4 and the Richardson No. (2-50m) was calculated based on a wind shear of 1.2 times the 2-80m shear. After fixing the bug, both calculations use the correct shears. -Afshin Andreas
- 12:43 MST: The BMS Direct TUVR QA Page has been updated to reflect a current problem we are having with the instrument. -Afshin Andreas
- 10:40 MST: The AOCS software was malfunctioning, the program was restarted today @ 10:39 MST. It is now working correctly. -Afshin Andreas
- 08:55 MST: The NWTC M2 Display software has been upgraded to a resolution of 1280x1024 for the (3) new LCD panels that will be installed at Bldg 251 and the IUF. Temperature has been added to the secondary axis of the plot and a time scale is now displayed on the x-axis. -Afshin Andreas
- 08:33 MST: Both batteries in the LI-1800's appear to be bad, both spectroradiometers stopped working on Feb 9 @ 8:20 MST, when we had a power failure. The instruments did not respond when power was restored. The Global 40-South LI-1800 was power cycled today at 8:30 MST and is functioning again. The Direct LI-1800 still does not work. -Afshin Andreas
- 13:47 MST: There is a problem with the NWTC M2 Temperature @ 50m, the signal is very noisy. -Afshin Andreas
- 12:49 MST: The Direct NIP at ECSU was replaced on January 18 at 16:10 EST, the new cal factor in the logger program was updated on January 30 at 14:40 EST. The CR10 temperature was also fixed (the internal CR10 temp is used instead of the CSI 107 temp sensor on CR10 wiring panel). The wind direction output was also upgraded, using the "wind vector" output option of the CR10 (before it was sampled every 5 minutes). -Afshin Andreas
- 11:13 MST: The NWTC M2 tower instruments were raised to their correct heights, today @ 10:21 MST. -Afshin Andreas
- 16:23 MST: The apparent problem with the NWTC M2 Wind Direction @ 80m has been corrected and the new aspirators fans are working properly. The tower instruments will be raised to their correct levels tomorrow. -Afshin Andreas
- 16:09 MST: Fixed a bug in the post-prosessing software, that tuncates the last digit from the raw output of the logger. This only applies to sites that have calculated fields added to the end of the data stream. -Afshin Andreas
- 08:46 MST: On January 25, 2002 at 8:00 MST, the NWTC M2 tower was dropped, for aspirator fan resplacement. The old baffle/microswitch system was being replaced with fans containing hall effect tachometers, using the datalogger we will now report fan speed in RPM. The Tower has not been raised due to an apparent problem with the wind direction at 80m. -Afshin Andreas
- 08:43 MST: On January 8, 2002 at 8:00 MST, the NWTC M2 tower was dropped, all wind speed & direction sensors and all temperature probes were replaced (except 2m temp). The temperature at 80m is now working again. The M2 Instrument page has been updated. -Afshin Andreas
- 09:58 MST: The Global TUVR cable was wired incorrectly yesterday. Corrected cable installed 0945 today. -Pete Gotseff
- 17:14 MST: The cable for the Global TUVR was replaced at 17:10 today.
-Pete Gotseff
- 11:14 MST: A BMS Instrument QA page was created for the Infrared CG4, which includes a statement about it being temporarily out of service. -Afshin Andreas
- 11:02 MST: A BMS Instrument QA page was created for the Global TUVR, which describes the current problem. -Afshin Andreas
- 11:13 MST: Fixed a bug when that occurs when performing a site comparision on two or more sites that have a data format change (or start) that occurs on the same day (RSP and RSP-Diagnostics for example). The bug would produce an extra invalid date range. -Afshin Andreas
- 08:05 MST: The outputs for the BMS CG4 Downwelling IR and Case Temp are now reading zero (to further clarify that this instrument is temporarly out of service). This was accomplished be disconnecting the Case Thermistor exciation voltage. -Afshin Andreas
- 16:21 MST: The BMS CG4 was removed from the diffuse tracker (to allow for another instrument to be calibrated), today at 11:25 MST. A shorting plug was put on the thermopile signal, net infrared is reading zero and downwelling infrared is reading garbage. -Afshin Andreas
- 09:41 MST: There is an intermitent problem with the BMS Global TUVR, some days its good and some days its bad. -Afshin Andreas
- 16:17 MST: There is still a problem with the BMS Global TUVR, signal is erratic. -Afshin Andreas
- 09:39 MST: An AOCS instrument description/history page has been created. -Afshin Andreas
- 09:08 MST: The BMS Global TUVR is working again, no problem was found. -Afshin Andreas
- 08:50 MST: There is a problem with the BMS Global TUVR, it's reading less than zero. (The problem started today). -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]